OpenStudy (anonymous):

SOMEONE PPLLLZZZ HELP!!!!!! solve (5/10x^2)=(1/10x)

OpenStudy (whpalmer4):

Is that \[\frac{5}{10x^2} = \frac{1}{10x}\]? Well, whatever it is, cross-multiplying should lead you to a solution. \[5*10x = 1*10x^2\]Cancel any common factors and solve.

OpenStudy (anonymous):

yes that is the problem

OpenStudy (anonymous):

so would the answer be 4=x

OpenStudy (anonymous):

x=0 or 1/5

OpenStudy (anonymous):

no 4 is not it

OpenStudy (anonymous):

but those ansers arent one of the choices

OpenStudy (anonymous):

A.5 B.1 C.-1 D.-5

OpenStudy (anonymous):

its either a or c

OpenStudy (anonymous):

A or B. Not C

OpenStudy (anonymous):

:( im prob gonna get it wrong

OpenStudy (anonymous):

i think its B.

OpenStudy (anonymous):

ok well thnx for ur help anyway, could u help me out with one more problem?

OpenStudy (anonymous):

you could use mathway its what ive been using most of High School.(:

OpenStudy (whpalmer4):

Come on, this is easy: after cross-multiplying, we have \[5∗10x=1∗10x^2\]\[50x=10x^2\]Do you see any common factors on both sides of the equals sign?

OpenStudy (whpalmer4):

If not, why don't you try dividing both sides of the equation by \(x\) and tell me what you get?